# Large Language Models for Computer-Aided Design: A Survey

### v1

@@ 6. Analysis and Discussion
OpenECAD (Yuan et al., 2024a) DeepCAD (Wu et al., 2021) Industry
@@
Text2CAD 1.0 & 1.1 (Khan et al., 2024b) DeepCAD Industry
Text-to-CAD (Wang et al., 2025b) DeepCAD, SkexGen (Xu et al., 2022b) Industry
CAD-GPT (Wang et al., 2025a) DeepCAD Industry
BlendNet (Du et al., 2024a) - Synthetic
LLM4CAD (Li et al., 2024b, 2025; Sun et al., 2025) - Synthetic
CADPrompt (Alrashedy et al., 2024) DeepCAD Industry
@@
CAD-Assistant (Mallis et al., 2024) SGPBench (Qiu et al., 2024) Industry
CAD-Recode (Rukhovich et al., 2024) - Synthetic
DeepCAD (Rukhovich et al., 2024; Zhang et al., 2025) - Industry
Fusion360 (Willis et al., 2021) (Rukhovich et al., 2024) - Industry
CC3D (Mallis et al., 2023) (Rukhovich et al., 2024) - Industry
QueryCAD (Kienle et al., 2024) ABC (Koch et al., 2019) Industry
SketchGraphs (Seff et al., 2020) (Wu et al., 2024) - Industry
Text2CAD (Li et al., 2024a) DeepCAD Industry
Img2CAD (You et al., 2024) ShapeNet (Chang et al., 2015) Synthetic
@@
CADTalk (Yuan et al., 2024b) - Synthetic & Industry
@@
Finally, we map each state-of-the-art work to its industrial context. This allows us to identify current active applications of LLMs for CAD, as well as the domains that are relatively underexplored. Table 5 presents the industries most commonly mentioned in the reviewed works. As shown in the table, the manufacturing industry currently attracts the most attention of the research community, whereas the shipbuilding industry has garnered relatively little focus. We also include the textile industry in the table which has no notable associated works. Nevertheless, we predict that future works will be associated to this industry due to the relevance of CAD to textile.
            Industry             Associated Works
            Automotive             (Yuan et al., 2024a), (Xu et al., 2025), (Du et al., 2024a), (Makatura et al., 2024), (Liu et al., 2023b),
            (Kodnongbua et al., 2023), (Wu et al., 2024), (Sun et al., 2024), (Picard et al., 2023), (Li et al., 2024e)
            Architecture             (Yuan et al., 2024a), (Xu et al., 2025), (Du et al., 2024a), (Makatura et al., 2023), (Makatura et al., 2024),
            (Wu et al., 2024), (You et al., 2024), (Zhang et al., 2025), (Sun et al., 2024), (Picard et al., 2023),
            (Li et al., 2024e)
            Shipbuilding             (Xu et al., 2025)
            Aerospace             (Yuan et al., 2024a), (Xu et al., 2025), (Du et al., 2024a), (Makatura et al., 2023), (Makatura et al., 2024),
            (Naik, 2024), (Yuan et al., 2024b), (Sun et al., 2024), (Picard et al., 2023), (Li et al., 2024e)
            Textile             -
            Manufacturing             (Yuan et al., 2024a), (Xu et al., 2025), (Khan et al., 2024b), (Wang et al., 2025b), (Wang et al., 2025a),
            (Du et al., 2024a), (Li et al., 2024a), (Li et al., 2024b), (Li et al., 2025), (Sun et al., 2025),
            (Alrashedy et al., 2024), (Badagabettu et al., 2024), (Makatura et al., 2023), (Makatura et al., 2024), (Nelson et al., 2023),
            (Mallis et al., 2024), (Rukhovich et al., 2024), (Jones et al., 2025), (Naik, 2024), (Kienle et al., 2024),n
            (Ocker et al., 2025), (Deng et al., 2024), (Wu et al., 2024), (You et al., 2024), (Zhang et al., 2025),
            (Tang et al., 2025), (Liu et al., 2023b), (Kodnongbua et al., 2023), (Yu et al., 2024), (Khan et al., 2024a),
            (Picard et al., 2023), (Wang et al., 2025c), (Li et al., 2024e)
            Medicine             (Makatura et al., 2023), (Makatura et al., 2024), (Sun et al., 2024), (Picard et al., 2023), (Li et al., 2024e)
            Electronics             (Makatura et al., 2023), (Makatura et al., 2024), (Sun et al., 2024), (Picard et al., 2023), (Li et al., 2024e)
            Consumer products             (Wu et al., 2024), (Zhang et al., 2025), (Liu et al., 2023b), (Yuan et al., 2024b), (Picard et al., 2023),
            (Wang et al., 2025c), (Li et al., 2024e)
            Entertainment             (Makatura et al., 2023), (Makatura et al., 2024)
            Education             (Picard et al., 2023), (Li et al., 2024e)
7. Future Directions
@@ 7.1. Application in Interior / Home Design
A promising future direction for LLMs in CAD is home design, also known as interior design. A few LLM-based tools for home design have already started to emerge (King et al., 2023; Çelen et al., 2024). For instance, HomeGPT [^5] https://www.homegpt.app/ focuses on redesigning individual rooms with AI, making it ideal for completing home makeovers or room-by-room transformations. RoomGPT [^6] https://www.roomgpt.io/ is another tool specializing in interior design, similar to ChatGPT, but tailored for creating and visualizing interior design styles and layouts. Users can upload a picture, and these systems generate a “dream space”. In (Çelen et al., 2024), I-Design utilized LLMs to transform text input into feasible scene graph designs, considering the relationships between objects. Littlefair et al. (Littlefair et al., 2025) demonstrated that LLMs could be combined with traditional optimization techniques to generate both functional and aesthetically pleasing interior designs. However, some of these tools remain closed source, and others do not focus on 3D CAD models. Consequently, home design represents a promising, but currently an underexplored area that is expected to attract more research attention in the future.
@@ 7.3. Building Compliance Checking in AEC
In the AEC industry, building designs must adhere to a wide range of requirements set forth by codes and standards (Ying and Sacks, 2024). Compliance with these regulations is essential to avoid legal issues, delays, and safety hazards (Saka et al., 2024). Traditionally, ensuring compliance has been a manual process, requiring labor-intensive checks using 2D drawings and documents, which is time-consuming and costly. Additionally, building codes are complex and frequently updated, further complicating the process. As a result, the need for Automatic Compliance Checking (ACC) has emerged as a critical solution. ACC involves two main steps: understanding design requirements from textual descriptions and analyzing 3D CAD models or 2D drawings to verify that they meet these requirements. Despite its importance, a universal and scalable ACC system remains a challenge (Amor and Dimyadi, 2021).
LLMs have already shown some promise in addressing this challenge within the AEC industry. For instance, Zheng and Fischer utilized GPT technologies to enhance natural language-based building information modeling (BIM) searches (Zheng and Fischer, 2023). Additionally, Ying and Sacks (Ying and Sacks, 2024) explored the use of LLMs for generic building compliance checking using BIM data. Du et al. (Du et al., 2024b) further leveraged LLMs to convert textual descriptions into code, generating editable BIM models and guiding iterative improvements in model quality. While these efforts focus on building infrastructure, there is a noticeable gap in research that directly applies LLMs to 3D CAD models. However, using LLMs to analyze 3D CAD models or 2D CAD drawings for automatic compliance checking holds significant potential, and further exploration in this area could greatly benefit the AEC industry.
@@ 7.4. Fashion AI and Textile Design
While much attention has been given to various applications of LLMs, fashion design, particularly in the context of textiles, remains largely unexplored. From Table 5, we observe that no state-of-the-art works have focused on leveraging LLMs for textile design. However, CAD plays a crucial role in fashion design, encompassing functions such as creative visualization, technical pattern development, and style information representation (Wu and Li, 2025).
CAD in fashion design has streamlined the exchange of information across complex network of communication channels, reducing time, cost, and material usage, while also improving accuracy and garment quality. This is largely due to the digitization of design information, which facilitates communication and decision-making without relying on physical samples. Given these advantages, fashion design presents an exciting opportunity for the application of LLMs in generating 3D CAD models. The potential to combine LLMs with fashion CAD tools could open up new avenues for creativity, efficiency, and innovation in the fashion industry.
@@ 8. Conclusion
In this comprehensive review, we explored the intersection of LLMs and CAD, highlighting the transformative potential of LLMs in various CAD-related tasks. We began by introducing the significance of CAD in industry, emphasizing its broad applications across different sectors. We then delved into the foundational principles of LLMs in an accessible manner, covering their architecture, training methods, and fine-tuning processes that enable them to perform across diverse domains. We also presented an overview of state-of-the-art LLMs, focusing on both closed-source models like GPT and PaLM and publicly available models such as LLaMA and DeepSeek. Our review underscored the remarkable capabilities of these models in handling complex tasks related to CAD, from data generation to coding and parametric generation, image synthesis, model evaluation and text generation. These advancements open up new possibilities for automating design workflows, improving efficiency, and enhancing creativity in CAD processes. Additionally, we critically analyzed current research trends, pinpointing emerging areas and potential gaps in the existing literature. Finally, we outlined promising future directions, including exploring LLMs in home design, specific data format generation, building compliance checking in AEC, and fashion design and textile applications. These areas offer exciting opportunities for future research and innovation, potentially leading to significant breakthroughs in the integration of LLMs with CAD.

### v2

@@ 8. Future Directions
OpenECAD (Yuan et al., 2024b) DeepCAD (Wu et al., 2021) Industry
@@
Text2CAD 1.0 & 1.1 (Khan et al., 2024a; Kolodiazhnyi et al., 2025; Chen et al., 2025) DeepCAD Industry
Text-to-CAD (Wang et al., 2025a) DeepCAD, SkexGen (Xu et al., 2022b) Industry
CAD-GPT (Wang et al., 2025b) DeepCAD Industry
BlendNet (Du et al., 2024) - Synthetic
LLM4CAD (Li et al., 2024c, 2025c; Sun et al., 2025) - Synthetic
CADPrompt (Alrashedy et al., 2024; Panta et al., 2025) DeepCAD Industry
@@
CAD-Assistant (Mallis et al., 2025) SGPBench (Qiu et al., 2025) Industry
CAD-Recode (Rukhovich et al., 2025; Kolodiazhnyi et al., 2025) - Synthetic
DeepCAD (Rukhovich et al., 2025; Zhang et al., 2025b, a) - Industry
Fusion360 (Willis et al., 2021) (Rukhovich et al., 2025; Kolodiazhnyi et al., 2025; Govindarajan et al., 2026) - Industry
CC3D (Mallis et al., 2023) (Rukhovich et al., 2025; Kolodiazhnyi et al., 2025) - Industry
QueryCAD (Kienle et al., 2025) ABC (Koch et al., 2019) Industry
SketchGraphs (Seff et al., 2020) (Wu et al., 2024b) - Industry
Text2CAD (Li et al., 2024d) DeepCAD Industry
Img2CAD (You et al., 2025) ShapeNet (Chang et al., 2015) Synthetic
@@
CADTalk (Yuan et al., 2024a; Chen et al., 2025) - Synthetic & Industry
@@
CAD-Llama (Li et al., 2025a) DeepCAD Synthetic & Industry
GenCAD-Code (Doris et al., 2025a) GenCAD (Alam and Ahmed, 2025) Synthetic
Text-to-CadQuery (Xie and Ju, 2025) DeepCAD, Text2CAD (Khan et al., 2024a) Industry
CAD-Editor (Yuan et al., 2025) DeepCAD Synthetic
CCC (Rosnitschek et al., 2025) - Synthetic
DesignQA (Doris et al., 2025b) - Industry
MPKG-WT (Hu et al., 2025) - Industry
Seek-CAD (Li et al., 2025d) ABC Industry
CAD2DMD (Valente et al., 2025) - Synthetic
DMDBench (Valente et al., 2025) - Industry
mrCAD (McCarthy et al., 2025) SketchGraphs Industry
CADmium (Govindarajan et al., 2026) DeepCAD Industry
DrivAerNet++(Elrefaie et al., 2024) (Elrefaie et al., 2025) - Industry
VideoCAD (Man et al., 2025) DeepCAD Synthetic
CAD-Coder (Guan et al., 2025) Text2CAD (Khan et al., 2024a) Industry
CADInstruct (Lv and Bao, 2025) DeepCAD Industry
TriView2CAD (Niu et al., 2025) - Synthetic & Industry
ShapeNet(Chang et al., 2015) (Yin et al., 2025) - Industry
CADReview (Chen et al., 2025) - Synthetic
@@ 8.1. Application in Interior / Home Design
A promising future direction for LLMs in CAD is home design, also known as interior design. A few LLM-based tools for home design have already started to emerge (King et al., 2023; Çelen et al., 2024). For instance, HomeGPT [^4] https://www.homegpt.app/ focuses on redesigning individual rooms with AI, making it ideal for completing home makeovers or room-by-room transformations. RoomGPT [^5] https://www.roomgpt.io/ is another tool specializing in interior design, similar to ChatGPT, but tailored for creating and visualizing interior design styles and layouts. Users can upload a picture, and these systems generate a “dream space”. In (Çelen et al., 2024), I-Design utilized LLMs to transform text input into feasible scene graph designs, considering the relationships between objects. Littlefair et al. (Littlefair et al., 2025) demonstrated that LLMs could be combined with traditional optimization techniques to generate both functional and aesthetically pleasing interior designs. However, some of these tools remain closed-source, and others do not focus on 3D CAD models. Consequently, home design represents a promising, but currently an underexplored area that is expected to attract more research attention in the future.
@@ 8.3. Building Compliance Checking in AEC
In the AEC industry, building designs must adhere to a wide range of requirements set forth by codes and standards (Ying and Sacks, 2024). Compliance with these regulations is essential to avoid legal issues, delays, and safety hazards (Saka et al., 2024). Traditionally, ensuring compliance has been a manual process, requiring labor-intensive checks using 2D drawings and documents, which is time-consuming and costly. Additionally, building codes are complex and frequently updated, further complicating the process. As a result, the need for Automatic Compliance Checking (ACC) has emerged as a critical solution. ACC involves two main steps: understanding design requirements from textual descriptions and analyzing 3D CAD models or 2D drawings to verify that they meet these requirements. Despite its importance, a universal and scalable ACC system remains a challenge (Amor and Dimyadi, 2021). LLMs have already shown some promise in addressing this challenge within the AEC industry. For instance, Zheng and Fischer utilized GPT technologies to enhance natural language-based building information modeling (BIM) searches (Zheng and Fischer, 2023). Additionally, Ying and Sacks (Ying and Sacks, 2024) explored the use of LLMs for generic building compliance checking using BIM data. Du et al. (Du et al., 2026) further leveraged LLMs to convert textual descriptions into code, generating editable BIM models and guiding iterative improvements in model quality. While these efforts focus on building infrastructure, there is a noticeable gap in research that directly applies LLMs to 3D CAD models. However, using LLMs to analyze 3D CAD models or 2D CAD drawings for automatic compliance checking holds significant potential, and further exploration in this area could greatly benefit the AEC industry.
Industry Automotive Architecture Shipbuilding Aerospace Textile Manufacturing Medicine Electronics Consumer products Entertainment Education
Associated Works (Yuan et al., 2024b), (Xu et al., 2025), (Du et al., 2024), (Makatura et al., 2024), (Liu et al., 2023b), (Kodnongbua et al., 2023), (Wu et al., 2024b), (Sun et al., 2024), (Picard et al., 2025), (Li et al., 2024e), (Govindarajan et al., 2026), (Elrefaie et al., 2025), (Man et al., 2025), (Lv and Bao, 2025) (Yuan et al., 2024b), (Xu et al., 2025), (Du et al., 2024), (Makatura et al., 2023), (Makatura et al., 2024), (Wu et al., 2024b), (You et al., 2025), (Zhang et al., 2025b), (Sun et al., 2024), (Picard et al., 2025), (Li et al., 2024e), (Xie and Ju, 2025), (Govindarajan et al., 2026), (Lv and Bao, 2025), (Madireddy et al., 2025), (Lee et al., 2026) (Xu et al., 2025) (Yuan et al., 2024b), (Xu et al., 2025), (Du et al., 2024), (Makatura et al., 2023), (Makatura et al., 2024), (Naik, 2024), (Yuan et al., 2024a), (Sun et al., 2024), (Picard et al., 2025), (Li et al., 2024e), (Govindarajan et al., 2026), (Man et al., 2025), (Lv and Bao, 2025) - (Yuan et al., 2024b), (Xu et al., 2025), (Khan et al., 2024a), (Wang et al., 2025a), (Wang et al., 2025b), (Du et al., 2024), (Li et al., 2024d), (Li et al., 2024c), (Li et al., 2025c), (Sun et al., 2025), (Alrashedy et al., 2024), (Badagabettu et al., 2024), (Makatura et al., 2023), (Makatura et al., 2024), (Nelson et al., 2023), (Mallis et al., 2025), (Rukhovich et al., 2025), (Jones et al., 2025), (Naik, 2024), (Kienle et al., 2025), (Ocker et al., 2025), (Deng et al., 2024), (Wu et al., 2024b), (You et al., 2025), (Zhang et al., 2025b), (Tang et al., 2025), (Liu et al., 2023b), (Kodnongbua et al., 2023), (Yu et al., 2024), (Khan et al., 2024b), (Picard et al., 2025), (Wang et al., 2025c), (Li et al., 2024e), (Li et al., 2025a), (Doris et al., 2025a), (Xie and Ju, 2025), (Kolodiazhnyi et al., 2025), (Yuan et al., 2025), (Li et al., 2025d), (Govindarajan et al., 2026), (Guan et al., 2025), (Lv and Bao, 2025), (Niu et al., 2025), (Zhang et al., 2025a), (Chen et al., 2025), (Lee et al., 2026), (Daareyni et al., 2025), (Ghosh et al., 2025), (Garcia et al., 2025), (Sadik and Bujny, 2025), (Rosnitschek et al., 2025), (Panta et al., 2025) (Makatura et al., 2023), (Makatura et al., 2024), (Sun et al., 2024), (Picard et al., 2025), (Li et al., 2024e) (Makatura et al., 2023), (Makatura et al., 2024), (Sun et al., 2024), (Picard et al., 2025), (Li et al., 2024e), (Valente et al., 2025) (Wu et al., 2024b), (Zhang et al., 2025b), (Liu et al., 2023b), (Yuan et al., 2024a), (Picard et al., 2025), (Wang et al., 2025c), (Li et al., 2024e), (Doris et al., 2025a), (Man et al., 2025), (Niu et al., 2025) (Makatura et al., 2023), (Makatura et al., 2024), (Govindarajan et al., 2026) (Picard et al., 2025), (Li et al., 2024e), (Zhang et al., 2024)
@@ 8.4. Fashion AI and Textile Design
While much attention has been given to various applications of LLMs, fashion design, particularly in the context of textiles, remains largely unexplored. From Table 5, we observe that no state-of-the-art works have focused on leveraging LLMs for textile design. However, CAD plays a crucial role in fashion design, encompassing functions such as creative visualization, technical pattern development, and style information representation (Wu and Li, 2025). CAD in fashion design has streamlined the exchange of information across complex network of communication channels, reducing time, cost, and material usage, while also improving accuracy and garment quality. This is largely due to the digitization of design information, which facilitates communication and decision-making without relying on physical samples. Given these advantages, fashion design presents an exciting opportunity for the application of LLMs in generating 3D CAD models. The potential to combine LLMs with fashion CAD tools could open up new avenues for creativity, efficiency, and innovation in the fashion industry.
8.5. Open Challenges
8.5.1. Multimodal Alignment
Currently, multimodal alignment across text, images, and parametric or code-based CAD models remains a crucial open challenge. Misalignment among these representations can lead to ambiguous reasoning, incorrect geometric interpretations, and unstable or unpredictable generation quality, ultimately limiting the reliability and controllability of LLM-driven CAD pipelines. In such misalignments, the first major type is image–instruction alignment, where visual content must accurately correspond to the textual descriptions of component positions, proportions, and specified conditions such as symmetry, rotation, and spatial relationships. (Kienle et al., 2025) leveraged ray casting (Glassner, 1989), together with additional post-processing steps, to perform image to CAD model alignment. The second major type is script–instruction alignment, where CAD scripts must faithfully implement attributes described in the instructions—including colors, sizes, materials, and other semantic properties that may not be directly observable in the images. Li et al. (Li et al., 2024d) aligned textual descriptions with parametric CAD sequences through a cascading contrastive strategy. In some cases, it is also necessary to align text, image, and point cloud data, in any combination. Hu et al. (Xu et al., 2025) proposed an image projection layer for aligning visual data and a point projection layer for aligning point cloud data. In (Wang et al., 2025b), a vision–language projector served as a bridge to align the two modalities.
8.5.2. Evaluation Inconsistency
From Section 6, we observe that different works adopt different evaluation metrics, making it difficult to compare methods fairly. Even for commonly used metrics such as Chamfer Distance, these global, non-differentiable measures provide no gradient information and map poorly to local geometric features. As a result, they offer limited guidance for refinement and can hinder the improvement steps performed by LLMs. We propose evaluating models using as many existing metrics as possible, and selecting a subset that best demonstrates model capabilities as a common and universal set of evaluation metrics.
8.5.3. Data Quality
From Table 4, we observe that most existing datasets are derived from DeepCAD, resulting in an overreliance on a single source and potentially constraining the generalizability of current research. Furthermore, a large portion of existing datasets are synthetic, raising concerns about data reliability and the extent to which models trained on such data can generalize to real-world scenarios. In addition, the majority of datasets are private, hindering reproducibility, limiting fair comparison across methods, and slowing down progress in the field. As such, there is a critical need for large-scale, diverse, and publicly accessible real-world datasets that capture a wide range of design settings, modeling workflows, and geometric complexities, thereby enabling more robust evaluation and fostering meaningful advances in LLM-assisted CAD research.
@@ 9. Conclusion
In this comprehensive review, we explored the intersection of LLMs and CAD, highlighting the transformative potential of LLMs in various CAD-related tasks. We began by introducing the significance of CAD in industry, emphasizing its broad applications across different sectors. We then delved into the foundational principles of LLMs in an accessible manner, covering their architecture, training methods, and fine-tuning processes that enable them to perform across diverse domains. We also presented an overview of state-of-the-art LLMs, focusing on both closed-source models like GPT and PaLM and publicly available models such as LLaMA and DeepSeek. Our review underscored the remarkable capabilities of these models in handling complex tasks related to CAD, from data generation to coding and parametric generation, image synthesis, model evaluation and text generation. These advancements open up new possibilities for automating design workflows, improving efficiency, and enhancing creativity in CAD processes. Importantly, we provide a comprehensive analysis of CAD evaluation, reviewing existing metrics and assessment methods in detail. Additionally, we critically analyzed current research trends, pinpointing emerging areas and potential gaps in the existing literature. Finally, we outlined promising future directions, including exploring LLMs in home design, specific data format generation, building compliance checking in AEC, and fashion design and textile applications. These areas offer exciting opportunities for future research and innovation, potentially leading to significant breakthroughs in the integration of LLMs with CAD.
10. Acknowledgments
This research was funded (partially) by the Australian Government through the Australian Research Council’s Discovery Early Career Researcher Award, project number DE220101057.
Dr. Naveed Akhtar is a recipient of the Australian Research Council Discovery Early Career Researcher Award (project number DE230101058) funded by the Australian Government.